The amount of CO2 in the atmosphere (as of 2019) is 412 ppm. Assuming perfect gas behavior, what is the partial pressure of CO2? ppm = parts per million, so the mole fraction of CO2 in air would be 412 x 10^-6
Answer #1

Mole fraction of CO2 in air = 412 * 10^-6

Or 4.12 * 10^-4

Atmospheric pressure = 1 atm ( total pressure of gas)

P(CO2) = mole fraction of CO2 * total pressure

= 4.12 * 10^-4 * 1

= 4.12 * 10^-4 atm

If you have any questions please comment

If you satisfied with the solution please rate it thanks